OpenStudy (anonymous):

x^3-5x^2-6x less than or equal to 0 answer is -1,6 need help solving.

OpenStudy (anonymous):

what do you need on

OpenStudy (anonymous):

help on

OpenStudy (anonymous):

I have the answers but don't know the process of solving it

OpenStudy (zzr0ck3r):

\(x^3-5x^2-6x\le 0\\x(x^2-5x-6)\le 0\\x(x+1)(x-6)\le0\) So we have roots at \(0,-1\) and \(6\). note that this thing has a positive coefficient on the first term, and it is a polynomial of degree three, so it looks like \(x^3\) for its end behavior, in other words it starts out negative. So since all of our roots are odd powers, it actually crosses the x-axis at each root (unlike \(x^2,x^4,...\)), so we start out negative, then cross the x-axis at \(x=-1\) and so we are positive now, then it crosses back over at \(x=0\) so now we are negative again, then it crosses back over the x-axis again at \(x=6\) then it remains positive from then on out (because there are no more roots). So finally to answer the question. The function is negative on \((-\infty, -1)\cup(0,6)\) Note that we use open brackets because it is NOT negative at the roots; it's \(0\) at the roots.

OpenStudy (zzr0ck3r):

woops, I forgot we were doing \(\le\) so it actually would be \[(-\infty, -1]\cup[0,6] \]

OpenStudy (campbell_st):

Just sketch it

OpenStudy (campbell_st):

And look at the parts below the x axis.... Then define the inequalities

OpenStudy (zzr0ck3r):

I agree! if you can use a calculator, or a graphing device. I normally work under the assumption that we cant use these things. You would need all the info I gave to sketch it without software?

OpenStudy (campbell_st):

well if you know the zeros and knowing the coefficient of the leading term is positive then the sketch is basically  now its easy to define the inequality without a lot of complicated work.